# Data Matters Most: Auditing Social Bias in Contrastive Vision–Language Models

### v5

@@ Abstract
Vision-language models (VLMs) deliver strong zero-shot recognition but frequently inherit social biases from their training data. We systematically disentangle three design factors—model size, training-data scale, and training-data source——by comparing CLIP and OpenCLIP, two models that share an identical contrastive objective yet differ in encoder width and in the image–text corpora on which they are pre-trained(400M proprietary pairs vs. 400M/2B LAION). Across balanced face-analysis benchmarks, enlarging the encoder reduces gender skew in CLIP but amplifies both gender and racial skew in OpenCLIP; increasing the LAION corpus from 400M to 2B further increases OpenCLIP bias. At matched model and data budgets, substituting proprietary data with LAION improves gender fairness while increasing racial skew, underscoring data source as the primary driver of bias patterns.
We also evaluate three post-hoc, test-time debiasing strategies — Bias Prompts, Prompt Array, and SANER. Debiasing reduces but does not eliminate harm, and its effectiveness is source- and size-dependent: Bias Prompts most effectively reduce gender skew in CLIP at smaller model sizes, whereas Prompt Array and SANER more reliably reduce racial skew in OpenCLIP; scaling LAION reconfigures which method is most fair. Taken together, these findings challenge the assumption that bigger models or datasets are automatically fairer and foreground training data source as the key determinant of both bias and mitigation efficacy. We release code and evaluation scripts to enable transparent, reproducible auditing of future VLMs. [^1] URL redacted for double-blind review
@@ 2 Related Work
We situate our study at the intersection of three longstanding research threads: (i) dataset bias in vision-language models (VLMs), (ii) bias-measurement and mitigation frameworks, and (iii) the scaling literature.
@@ 2.1 Dataset Bias in Vision–Language Models
Early audits showed that contrastive VLMs reproduce demographic co-occurrence statistics present in web-scale pre-training data. Agarwal et al. (2021) uncovered disproportionate associations between racial descriptors and CLIP embeddings, while Birhane et al. (2021) documented misogynistic and racist imagery in LAION-400M. Subsequent benchmarks broadened the evidence base: ModSCAN probes gender and race stereotypes across occupations and persona traits (Jiang et al., 2024), So-B-IT demonstrates that toxic prompts such as terrorist yield demographically skewed retrievals (Hamidieh et al., 2024), and VisoGender targets gender resolution bias in image–text pronoun disambiguation (Hall et al., 2023). Very recent work highlights cultural and socioeconomic blind spots introduced by English-only filtering of web data (Pouget et al., 2024).
@@ 2.2 Bias-Measurement Frameworks and Mitigation
Task-agnostic toolkits now quantify social bias in VLM outputs. MMBias measures stereotype leakage in captioning and VQA (Janghorbani & de Melo, 2023), while DebiasCLIP reduces gender stereotypes by prepending learned visual tokens (Berg et al., 2022b). Embedding-space analyses further show that CLIP encodes the family $\leftrightarrow$ career gender stereotype (Bianchi et al., 2023; Liang et al., 2021). Prompt-level interventions—e.g., the Debiasing and VisDebiasing prefixes in ModSCAN—can attenuate skew but rarely eliminate it (Jiang et al., 2024).
@@ 2.3 Scaling Effects on Bias
Scaling model parameters or data volume is often viewed as a route to robustness, yet empirical findings remain mixed. In language models, larger GPT-style systems sometimes reduce toxic generation (Hoffmann et al., 2022) but can entrench occupational stereotypes (Liang et al., 2021). Ghate et al. (2025) shows that intrinsic bias is largely predictable from pre-training data, not architecture. Liu et al. (2024) offers a probabilistic framework revealing that stereotype volatility increases with model size. For VLMs, Birhane et al. (2023) warned that expanding LAION from 400M to 5B images risks amplifying existing harms, whereas higher image resolution can mitigate certain gender biases. Our controlled study extends this line by showing that size, scale, and loss source interact non-linearly: the same parameter increase that softens bias in CLIP amplifies it in OpenCLIP, and a five-fold data increase leaves CLIP’s skew almost unchanged while doubling that of OpenCLIP.
@@ 3.3 Bias Probing Tasks
Crime vs. Non human.
Following the original CLIP paper (Radford et al., 2021), we contrast ‘‘photo of a <label> criminal’’ with ‘‘photo of a <label> animal’’, where <label> instantiates the protected attribute under study (e.g., “Black man”, “Asian woman”). The template covers gender, race, and a Non human catch-all to test whether models associate certain demographics more strongly with criminality.
@@ 3.6 Inference Details
Images are resized to $224\times 224$ for ViT B/32 and $336\times 336$ for ViT L/14 to match pre-training. Caption embeddings use the checkpoint-specific temperature $\tau$ without test time augmentation. BP uses the authors’ original prompt pairs and the calibrated projection matrix released with the paper. PA is trained for three epochs on 90k FairFace images plus the same number of LAION images, using the hyperparameters from the AACL-22 code release ($\lambda_{\text{ITC}}{=}0.05$). SANER follows the SD-XL caption protocol of Hirota et al. (2025) and is trained for five epochs on COCO 2017. During inference, the debias layer or token array is inserted after all pre-processing so that the downstream evaluation pipeline (§LABEL:ssec:_tasks) remains identical. All three methods operate without retraining the backbone encoders. Bias Prompts and SANER adjust text features at inference (closed form or light layer), while Prompt Array learns a small set of tokens with an adversary to reduce attribute leakage. A single NVIDIA A100 (40 GB) processes the full benchmark in under thirty minutes. Code, prompts, and raw outputs will be released upon publication.
@@ 5 Discussion
Our experiments overturn the convenient intuition that higher capacity inevitably brings greater fairness. When CLIP grows from ViT-B/32 to ViT-L/14, gender skew falls by roughly one-third while race skew stays level; the identical size increase inside OpenCLIP instead amplifies both forms of skew. Because the loss function and optimisation recipe are shared, the divergent outcomes must arise from how each family’s pre-training corpus shapes the gradients that extra parameters can exploit. In other words, parameter count is only as fair as the data distribution that guides it.
@@ Appendix A Neutral-Image Control
Appendix B Calibration of Directional-Bias Magnitude
@@
We calibrate directional bias magnitudes to practical risk by mapping the absolute cosine difference $|\Delta|$ to the probability that the model outputs a harmful label: the negative-pole prompt for Comm/Agen, or an animal/criminal term for Crime.
@@
A flat segment near $p_{\mathrm{harm}}\!\approx\!0.1$ indicates chance-level behavior. The elbow is the first bin with $p_{\mathrm{harm}}\!\geq\!0.5$; the plateau indicates saturation. Figure 10 shows a typical shape.
@@
(i) Bias magnitudes below $0.1$ are effectively benign: $|\Delta|<0.1$ shifts $p_{\mathrm{harm}}$ by <2<\!2 pp, matching neutral-image controls. (ii) For CLIP ViT-B/32, risk crosses $50\%$ around $|\Delta|\!\approx\!0.25$ and reaches $\geq\!90\%$ beyond $0.5$, marking moderate vs. severe bias thresholds. (iii) OpenCLIP curves are steeper: elbows often appear in the 0.1–0.2 bin and $p_{\mathrm{harm}}\!\geq\!95\%$ by $|\Delta|=0.3$, consistent with the larger max-skew values in the main text. (iv) Crime saturates fastest, indicating that non-human or criminal mislabellings are particularly sensitive to directional bias.
These calibration curves turn cosine magnitudes into actionable thresholds: a seemingly small $|\Delta|=0.3$ already implies a four-in-five chance of a harmful prediction for several checkpoints, making directional bias a practically meaningful metric.
Appendix C Robustness to Prompt-Template Variation

### v6

@@ Abstract
Vision-language models (VLMs) deliver strong zero-shot recognition but frequently inherit social biases from their training data. We systematically disentangle three design factors—model size, training‑data scale, and training‑data source——by comparing CLIP and OpenCLIP, two models that share an identical contrastive objective yet differ in encoder width and in the image–text corpora on which they are pre‑trained(400M proprietary pairs vs. 400M/2B LAION). Across balanced face-analysis benchmarks, enlarging the encoder reduces gender skew in CLIP but amplifies both gender and racial skew in OpenCLIP; increasing the LAION corpus from 400M to 2B further increases OpenCLIP bias. At matched model and data budgets, substituting proprietary data with LAION improves gender fairness while increasing racial skew, underscoring data source as the primary driver of bias patterns.
We also evaluate three post-hoc, test-time debiasing strategies — Bias Prompts, Prompt Array, and SANER. Debiasing reduces but does not eliminate harm, and its effectiveness is source- and size-dependent: Bias Prompts most effectively reduce gender skew in CLIP at smaller model sizes, whereas Prompt Array and SANER more reliably reduce racial skew in OpenCLIP; scaling LAION reconfigures which method is most fair. Taken together, these findings challenge the assumption that bigger models or datasets are automatically fairer and foreground training data source as the key determinant of both bias and mitigation efficacy. We release code and evaluation scripts to enable transparent, reproducible auditing of future VLMs. [^1] Code available at https://github.com/zahraaalsahili/CLIP_Bias.
@@ 2 Related Work
We situate our study at the intersection of three longstanding research threads: (i) dataset bias in vision-language models (VLMs), (ii) bias‑measurement and mitigation frameworks, and (iii) the scaling literature.
@@ 2.1 Dataset Bias in Vision–Language Models
Early audits showed that contrastive VLMs reproduce demographic co‑occurrence statistics present in web‑scale pre‑training data. Agarwal et al. (2021) uncovered disproportionate associations between racial descriptors and CLIP embeddings, while Birhane et al. (2021) documented misogynistic and racist imagery in LAION‑400M. Subsequent benchmarks broadened the evidence base: ModSCAN probes gender and race stereotypes across occupations and persona traits (Jiang et al., 2024), So‑B‑IT demonstrates that toxic prompts such as terrorist yield demographically skewed retrievals (Hamidieh et al., 2024), and VisoGender targets gender resolution bias in image–text pronoun disambiguation (Hall et al., 2023). Very recent work highlights cultural and socioeconomic blind spots introduced by English‑only filtering of web data (Pouget et al., 2024).
@@ 2.2 Bias-Measurement Frameworks and Mitigation
Task-agnostic toolkits now quantify social bias in VLM outputs. MMBias measures stereotype leakage in captioning and VQA (Janghorbani & deăMelo, 2023), while DebiasCLIP reduces gender stereotypes by prepending learned visual tokens (Berg et al., 2022b). Embedding-space analyses further show that CLIP encodes the family $\leftrightarrow$ career gender stereotype (Bianchi et al., 2023; Liang et al., 2021). Prompt-level interventions—e.g., the Debiasing and VisDebiasing prefixes in ModSCAN—can attenuate skew but rarely eliminate it (Jiang et al., 2024).
@@ 2.3 Scaling Effects on Bias
Scaling model parameters or data volume is often viewed as a route to robustness, yet empirical findings remain mixed. In language models, larger GPT‑style systems sometimes reduce toxic generation (Hoffmann et al., 2022) but can entrench occupational stereotypes (Liang et al., 2021). Ghate et al. (2025) shows that intrinsic bias is largely predictable from pre‑training data, not architecture. Liu et al. (2024) offers a probabilistic framework revealing that stereotype volatility increases with model size. For VLMs, Birhane et al. (2023) warned that expanding LAION from 400M to 5B images risks amplifying existing harms, whereas higher image resolution can mitigate certain gender biases. Our controlled study extends this line by showing that size, scale, and loss source interact non‑linearly: the same parameter increase that softens bias in CLIP amplifies it in OpenCLIP, and a five‑fold data increase leaves CLIP’s skew almost unchanged while doubling that of OpenCLIP.
@@ 3.3 Bias Probing Tasks
Crime vs. Non-Human.
Following CLIP-style probing (Radford et al., 2021), we test whether images of different demographics align more with crime-related or non-human concepts. For each image $I$ with demographic label $d$, we compute CLIP similarities to all prompts in the Crime-Related set and, separately, the Non-Human set, pool scores within each set (max unless stated), and compare the pooled scores to place $I$ on the axis. This probe tests whether certain demographics are more associated with crime concepts than non-human concepts without text concatenation.
@@ 3.6 Inference Details
Images are resized to $224\times 224$ for ViT B/32 and $336\times 336$ for ViT L/14 to match pre-training. Caption embeddings use the checkpoint-specific temperature $\tau$ without test time augmentation. BP uses the authors’ original prompt pairs and the calibrated projection matrix released with the paper. PA is trained for three epochs on 90k FairFace images plus the same number of LAION images, using the hyperparameters from the AACL-22 code release ($\lambda_{\text{ITC}}{=}0.05$). SANER follows the SD-XL caption protocol of Hirota et al. (2025) and is trained for five epochs on COCO 2017. During inference, the debias layer or token array is inserted after all pre-processing so that the downstream evaluation pipeline remains identical. All three methods operate without retraining the backbone encoders. Bias Prompts and SANER adjust text features at inference (closed form or light layer), while Prompt Array learns a small set of tokens with an adversary to reduce attribute leakage. A single NVIDIA A100 (40 GB) processes the full benchmark in under thirty minutes. Code, prompts, and raw outputs will be released upon publication.
@@ 5 Discussion
Our experiments overturn the convenient intuition that higher capacity inevitably brings greater fairness. When CLIP grows from ViT‐B/32 to ViT‐L/14, gender skew falls by roughly one-third while race skew stays level; the identical size increase inside OpenCLIP instead amplifies both forms of skew. Because the loss function and optimisation recipe are shared, the divergent outcomes must arise from how each family’s pre-training corpus shapes the gradients that extra parameters can exploit. In other words, parameter count is only as fair as the data distribution that guides it.
@@ 6 Conclusion
Acknowledgements
MP was supported by the UK Engineering and Physical Sciences Research Council via Responsible AI UK (grant number EP/Y009800/1, project KP0016, “AdSoLve: Addressing Socio-technical Limitations of LLMs for Medical and Social Computing”); by the European Union’s Horizon Europe research and innovation programme under grant agreement number 101214398 (ELLIOT); and by the Slovenian Research and Innovation Agency (ARIS) through the Gravitacije project LLM4DH (“Large Language Models for Digital Humanities”, GC-0002), the project CroDeCo (“Cross-Lingual Analysis for Detection of Cognitive Impairment in Less-Resourced Languages”, J6-60109) and the research programme “Knowledge Technologies” (P2-0103). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. ZA is supported by Google DeepMind PhD Fellowship and thanks their Google DeepMind mentor, David Stutz, for guidance and support.
@@ Appendix A Neutral-Image Control
Appendix B Calibration of Directional‐Bias Magnitude
@@
We calibrate directional bias magnitudes to practical risk by mapping the absolute cosine difference $|\Delta|$ to the probability that the model outputs a harmful label: the negative‐pole prompt for Comm/Agen, or an animal/criminal term for Crime.
@@
A flat segment near $p_{\mathrm{harm}}\!\approx\!0.1$ indicates chance‐level behavior. The elbow is the first bin with $p_{\mathrm{harm}}\!\geq\!0.5$; the plateau indicates saturation. Figure 10 shows a typical shape.
@@
(i) Bias magnitudes below $0.1$ are effectively benign: $|\Delta|<0.1$ shifts $p_{\mathrm{harm}}$ by <2<\!2 pp, matching neutral‐image controls. (ii) For CLIP ViT‐B/32, risk crosses $50\%$ around $|\Delta|\!\approx\!0.25$ and reaches $\geq\!90\%$ beyond $0.5$, marking moderate vs. severe bias thresholds. (iii) OpenCLIP curves are steeper: elbows often appear in the 0.1–0.2 bin and $p_{\mathrm{harm}}\!\geq\!95\%$ by $|\Delta|=0.3$, consistent with the larger max‐skew values in the main text. (iv) Crime saturates fastest, indicating that non‐human or criminal mislabellings are particularly sensitive to directional bias.
These calibration curves turn cosine magnitudes into actionable thresholds: a seemingly small $|\Delta|=0.3$ already implies a four‐in‐five chance of a harmful prediction for several checkpoints, making directional bias a practically meaningful metric.
Appendix C Robustness to Prompt‐Template Variation